# New perspectives on the wave equation for a massive scalar field in Schwarzschild spacetime

### v1

@@ Appendix B Canonical reduction with respect to $z$
It is very easy to perform the canonical reduction of the radial equation (31) with respect to the independent variable $z$. Following the same procedure of Sec. II, we obtain
@@
The function $\beta_{l\omega}$ is the solution of the differential equation
@@
In the neighbourhood of $z=0$, the potential takes the approximate form
@@
In the neighbourhood of $z=1$ we find instead (after imposing regularity of $\beta$ therein)
@@
We note complete correspondence with the leading behaviours obtained in Sec. III with respect to the $r$ variable.
@@ Appendix C Remarks on the solution at infinity
First, we provide a theoretical derivation of the ansatz given by Eq. (42) based on a method developed by Poincaré Poincare ; Poincare2 to compute the solution of a differential equation with variable coefficients. We then evaluate the explicit form of the first coefficients $B_{k}$ of the solution given by Eq. (49).
@@ C.1 Theoretical derivation
The radial wave equation in its normal form (9), with the potential term defined in Eq. (10), can be solved locally, i.e., in the neighbourhhod of the singular points Persides1 ; Fiziev ; Philipp ; Minucci . For this purpose, it is helpful to re-express it in a form where the variable coefficients are polynomials, following the seminal work of Poincaré Poincare . This is achieved by multiplying it on the left by $r^{2}(r-2M)^{2}$. This leads to the differential equation
@@
$P_{0l\omega\mu}(r)=r^{2}(r-2M)^{2}J_{l\omega\mu}(r)$ (99)
$=$ $(\omega^{2}-\mu^{2})r^{4}+2M\mu^{2}r^{3}+\left[\frac{1}{2}-\frac{1}{2}-l(l+1)\right]r^{2}$
$+$ $\Bigr{[}-M+M+2Ml(l+1)\Bigr{]}r+M^{2}$
$=$ $(\omega^{2}-\mu^{2})r^{4}+2M\mu^{2}r^{3}-l(l+1)r^{2}+2Ml(l+1)r+M^{2},$
and we have stressed in the intermediate step the exact cancellations occurring in the evaluation of the polynomial (99). As far as we know, only the authors of Ref. Philipp came pretty close to studying an equation like our (97) with $r$ as independent variable, but they chose a third degree polynomial in their Eq. (32) as the coefficient of the second derivative of the desired solution.
At this stage, the approximate solutions (17), (27) and the one which would be obtained for Eq. (28) by considering only the leading term (labelled as $\beta_{\omega,\infty}$) suggest looking for a solution of Eq. (97) in the factorized form of product (called $s$-homotopic ansatz in the current literature Philipp ; Minucci )
@@
where the function $G_{l\omega\mu}$ should be such that this product solves, in the neighbourhood of $r=0,2M,\infty$, Eq. (97). In the ansatz (100) we have
$\beta_{0}(r)\beta_{\omega,2M}(r)\beta_{\omega,\infty}(r)=\sqrt{r(r-2M)}\Bigr{[}D_{1}(r-2M)^{2iM\omega}+D_{2}(r-2M)^{-2iM\omega}\Bigr{]}$ (101)
$\times$ $\Bigr{[}S_{1}e^{r\sqrt{\mu^{2}-\omega^{2}}}+S_{2}e^{-r\sqrt{\mu^{2}-\omega^{2}}}\Bigr{]}.$
The factorization in Eq. (100) follows what was learned from Poincaré Poincare2 , i.e., for linear differential equations with variable coefficients, the solution consists of the limiting form at the origin times the limiting form at infinity times a suitable series. With respect to the case considered in Ref. Poincare2 , we are here dealing with two (rather than just one) regular singular points at finite values of the independent variable. The structure of the ansatz given by Eqs.(100), (101) is too rich and it does not provide any semplification of computations. Our proposal, given by Eq. (42), is an easier case where $S_{1}=0$, $\beta_{0}(r)\beta_{\omega,2M}(r)\simeq r^{\alpha}$ and the interpolating series $F_{l\omega}(r)$ consists only of negative powers of r. In this way we are able to study the behaviour of the field at infinity without any effort.
@@ C.2 Some coefficients
We begin by writing the first coefficients $A_{k}$ of the asymptotic expansion of the canonical potential at infinity, given by Eq. (28)
@@
In ligth of the above, the first two $B_{s}$ coefficients are explicitly obtained here:
@@ Appendix D Comparison with the Persides analysis
Sections VI and VII have presented our original calculations for a massive scalar field in a fixed Schwarzschild background, whereas, for a massless field, the large-$r$ solution of the radial wave equation was first obtained by Persides Persides1 by using a $3$-term recurrence relation, because he did not reduce the equation for Fourier modes $R_{l\omega}$ to its normal form. With our notation, the Persides large-$r$ form of $R_{l\omega}(r)$ reads as (setting $c$ and $G$ to $1$ in Ref. Persides1 )
@@
On the other hand, in light of our Sec. VI, one might be tempted to study the massless case by defining
@@
However, such a formula cannot be taken to hold, because Eq. (49) is only a Poincaré asymptotic expansion (see Appendix E), and hence
@@
As far as we can see, the best that one can do is to check that, by including as many $B_{s}(l,\omega,\mu)$ coefficients as possible in the evaluation of the massive case, and as many $\tau_{n}(l,\omega)$ coefficients as possible in the massless case, one obtains an accurate approximation of the local solution of the radial wave equation in the neighbourhood of $r=\infty$. In the latter case, one should also bear in mind that Persides1
@@ Appendix E The Poincaré framework
In Ref. Poincare2 , Poincaré studied linear differential equations of the form
@@
which satisfy formally Eq. (116), where $Q(x)$ is a polynomial in $x$. The series in Eq. (117) was then said to be a normal series. In particular, if the normal series is convergent, one says that Eq. (116) admits a normal integral.
@@
where the sum of the first $(n+1)$ terms is denoted by $S_{n}$, represents asymptotically a function $J:x\rightarrow J(x)$ if the epression
@@
This means that, for $x$ suffiently large, one has the majorization
@@
The literature on theoretical physics still relies upon the Poincaré definition of asymptotic expansion, whereas in pure mathematics one has also a different, rigorous framework built by Dieudonné Dieudonne .

### v2

@@ Appendix B Canonical Reduction with Respect to $z$
It is very easy to perform the canonical reduction of the radial equation (9) with respect to the independent variable $z$. Following the same procedure of Section II, we obtain
@@
The function $\beta_{l\omega\mu}$ is the solution of the differential equation
@@
In the neighborhood of $z=0$, the potential takes the approximate form
@@
In the neighborhood of $z=1$ we find instead (after imposing regularity of $\beta$ therein)
@@
We note complete correspondence with the leading behaviors obtained in Section III with respect to the $r$ variable.
@@ Appendix C Remarks on the Solution at Infinity
First, we provide a theoretical derivation of the ansatz given by Equation (42) based on a method developed by Poincaré [22, 23] to compute the solution of a differential equation with variable coefficients. We then evaluate the explicit form of the first coefficients $B_{k}$ of the solution given by Equation (49).
@@ C.1 Theoretical Derivation
The radial wave equation in its normal form (9), with the potential term defined in Equation (10), can be solved locally, i.e., in the neighborhood of the singularpoints [3, 17, 18, 21]. For this purpose, it is helpful to re-express it in a form where the variable coefficients are polynomials, following the seminal work of Poincaré [22]. This is achieved by multiplying it on the left by $r^{2}(r-2M)^{2}$. This leads to the differential equation
@@
$\begin{array}[]{lll}&&P_{0l\omega\mu}(r)=r^{2}(r-2M)^{2}J_{l\omega\mu}(r)\\ &=&(\omega^{2}-\mu^{2})r^{4}+2M\mu^{2}r^{3}+\left[\frac{1}{2}-\frac{1}{2}-l(l+1)\right]r^{2}\\ &+&\Bigr{[}-M+M+2Ml(l+1)\Bigr{]}r+M^{2}\\ &=&(\omega^{2}-\mu^{2})r^{4}+2M\mu^{2}r^{3}-l(l+1)r^{2}+2Ml(l+1)r+M^{2},\end{array}$ (99)
and we have stressed in the intermediate step the exact cancellations occurring in the evaluation of the polynomial (99). As far as we know, only the authors of Ref. [18] came pretty close to studying an equation like our (97) with $r$ as an independent variable, but they chose a third degree polynomial in their Equation (32) as the coefficient of the second derivative of the desired solution.
At this stage, the approximate solutions (17), (27) and the one which would be obtained for Equation (28) by considering only the leading term (labelled as $\beta_{\omega,\infty}$) suggest looking for a solution of Equation (97) in the factorized form of product (called $s$-homotopic ansatz in the current literature [18, 21])
@@
where the function $G_{l\omega\mu}$ should be such that this product solves, in the neighborhood of $r=0,2M,\infty$, Equation (97). In the ansatz (100) we have
$\begin{array}[]{lll}&&\beta_{0}(r)\beta_{\omega,2M}(r)\beta_{\omega,\infty}(r)=\sqrt{r(r-2M)}\Bigr{[}D_{1}(r-2M)^{2iM\omega}+D_{2}(r-2M)^{-2iM\omega}\Bigr{]}\\ &\times&\Bigr{[}S_{1}e^{r\sqrt{\mu^{2}-\omega^{2}}}+S_{2}e^{-r\sqrt{\mu^{2}-\omega^{2}}}\Bigr{]}.\end{array}$ (101)
The factorization in Equation (100) follows what was learned from Poincaré [23], i.e., for linear differential equations with variable coefficients, the solution consists of the limiting form at the origin times the limiting form at infinity times a suitable series. With respect to the case considered in Ref. [23], we are here dealing with two (rather than just one) regular singular points at finite values of the independent variable. The structure of the ansatz given by Equations (100), (101) is too rich, and it does not provide any simplification of computations. Our proposal, given by Equation (42), is an easier case where $S_{1}=0$, $\beta_{0}(r)\beta_{\omega,2M}(r)\simeq r^{\alpha}$ and the interpolating series $G_{l\omega\mu}(r)$ consists only of negative powers of r. In this way we are able to study the behavior of the field at infinity without any effort.
@@ C.2 Some Coefficients
We begin by writing the first coefficients $A_{k}$ of the asymptotic expansion of the canonical potential at infinity, given by Equation (28):
@@
In light of the above, the first two $B_{s}$ coefficients are explicitly obtained here:
@@ Appendix D Comparison with the Persides Analysis
Sections VI and VII have presented our original calculations for a massive scalar field in a fixed Schwarzschild background, whereas, for a massless field, the large-$r$ solution of the radial wave equation was first obtained by Persides [3] by using a $3$-term recurrence relation, because he did not reduce the equation for Fourier modes $R_{l\omega}$ to its normal form. With our notation, the Persides large-$r$ form of $R_{l\omega}(r)$ reads as (setting $c$ and $G$ to $1$ in Ref. [3])
@@
On the other hand, in light of our Section VI, one might be tempted to study the massless case by defining
@@
However, such a formula cannot be taken to hold, because Equation (49) is only a Poincaré asymptotic expansion (see Appendix E [16]), and hence
@@
As far as we can see, the best that one can do is to check that by including as many $B_{s}(l,\omega,\mu)$ coefficients as possible in the evaluation of the massive case, and as many $\tau_{n}(l,\omega)$ coefficients as possible in the massless case, one obtains an accurate approximation of the local solution of the radial wave equation in the neighborhood of $r=\infty$. In the latter case, one should also bear in mind that [3]
@@ Appendix E The Poincaré Framework
In Ref. [23], Poincaré studied linear differential equations of the form
@@
which satisfy formally Equation (116), where $Q(x)$ is a polynomial in $x$. The series in Equation (117) was then said to be a normal series. In particular, if the normal series is convergent, one says that Equation (116) admits a normal integral.
@@
where the sum of the first $(n+1)$ terms is denoted by $S_{n}$, represents asymptotically a function $J:x\rightarrow J(x)$ if the expression
@@
This means that, for $x$ sufficiently large, one has the majorization
@@
The literature on theoretical physics still relies upon the Poincaré definition of asymptotic expansion, whereas in pure mathematics one has also a different, rigorous framework built by Dieudonné [30].